OpenStudy (anonymous):

A rectangle has length of x+9 and width 2x-1. What is the area of the rectangle?

OpenStudy (unklerhaukus):

The area of a rectangle is the length times the width \[\text{Area}=\text{length}\times \text{width}\]\[A=l\times w\]\[A=(x+9)\times(2x-1)\]

OpenStudy (unklerhaukus):

\[(x+9)\times(2x−1)=2x^2-x+18x-9=2x^2+17x-9\]